Rearrange Numbers in String


Given a string, write a program to re-arrange all the numbers appearing in the string in decreasing order. Note: There will not be any negative numbers or numbers with decimal part.


Input


The input will be a single line containing a string.


Output


The output should be a single line containing the modified string with all the numbers in string re-ordered in decreasing order.


Explanation


For example, if the given string is "I am 5 years and 11 months old", the numbers are 5, 11. Your code should print the sentence after re-ordering the numbers as "I am 11 years and 5 months old".


Sample Input

I am 5 years and 11 months old

Sample Output

I am 11 years and 5 months old

Sample Input 2:

I am 11 years 5 months, 26 weeks and 38 days old

Sample output:

I am 38 years 26 months, 11 weeks and 5 days old


1
Expert's answer
2021-07-12T03:09:30-0400
print("Enter a string")
s = input()


li = list(s.split(" "))


li = [int(i) for i in li]


li.sort(reverse = True)


list_string = map(str, li)


s = ""
for x in list_string:
    s = s + x + " "
    
print(s)